OpenStudy (anonymous):

n3 + 2n is divisible by 3 for all n≥ 1 by induction

myininaya (myininaya):

So we want to show 3|(n^3+2n) for all n>=1. You basically have two steps, but I will write 3. 1) Show it is true for n=1. 2) Assume it is true for some integer k greater than 1. (no work to be done on step 2) 3) Show it is true for k+1 using step 2.

OpenStudy (anonymous):

thank u

myininaya (myininaya):

Let me know if you run into trouble with the steps.

OpenStudy (anonymous):

i need the answer for my assignment so i can use what u wrote ?

myininaya (myininaya):

For 1 and 3 you have to do some work.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank u once again

myininaya (myininaya):

So is that n^3+2n? Did I translate that right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

I find it helpful to recall that 3|(n^3+2n) just means we have 3m=n^3+2n for some integer m. Whenever one of my proofs that call for induction has that little divides symbol I use that definition. If a|b then ak=b for integer k.

OpenStudy (zarkon):

I like this ... \[n^3+2n=n^3-n+3n=n(n^2-1)+3n=n(n-1)(n+1)+3n\] \[=(n-1)n(n+1)+3n\]

OpenStudy (zarkon):

no induction required ;)

OpenStudy (zarkon):

since \(3|(n-1)n(n+1)\) and \(3|3n\)

myininaya (myininaya):

I'm trying to think why (n-1)n(n+1) would be divisible by 3. I mean of course I could go through numbers to see it is divisible by 3 but that would be just induction again. What elementary thingy am I missing?

OpenStudy (zarkon):

it is the product of 3 consecutive numbers (mod 3 one of them must be 0)

myininaya (myininaya):

also it is every 3 integer numbers. 1 2 3 2 3 4 4 5 6 one of those numbers has to be divsible by 3 Okay.

myininaya (myininaya):

Yea, consecutive. Whatever. Nice Zarkon.

OpenStudy (zarkon):

though I'm sure the OP is learning induction and this approach would get them 0 points :)